OpenStudy (calculusxy):

MEDAL FOR THE BEST ANSWER! Rodney has fifty coins, including at least one quarter, that are worth $1.00. If he loses one coin, what is the probability that it was a dime? A. 0.01 B. 0.02 C. 0.04 D. 0.05 E. 0.10 @Mertsj

OpenStudy (anonymous):

e cause it says that only one of them is a quarter

OpenStudy (calculusxy):

But that is not correct according to the answer key.

OpenStudy (anonymous):

then...

OpenStudy (mertsj):

So $1 = 1 quarter + 2 dimes + 2 nickels + 45 pennies P(dime)=2/50=1/25=.04

OpenStudy (calculusxy):

What do you mean by P?

OpenStudy (charlotte123):

P - Probability. @calculusxy

OpenStudy (calculusxy):

How about where did he get 2/50 from?